# Nuclear Astrophysics

### v1

@@ 3.2.1 Single stars
As a result of these various fading central energy sources towards the end of stellar evolution of stars with masses exceeding $\simeq$ 8 M, gravitational collapse will take over, and proceed faster with electron capture and hence neutronization increasing in central regions [see 231, for a recent review]. This neutronization of matter near and in the newly-forming neutron star in the center of gravitational collapse results in an additional and very intense burst of electron neutrinos. This is one of the significant messengers for astronomical measurement (see section 4 on Nuclear Astronomy).
The collapse is halted when central density reaches values above nuclear saturation density beyond 2.7 1014g cm-3, as compression then faces the repulsive part of the strong nuclear potential. This is reflected in the equation of state of matter in the newly-forming neutron star [229], which describes the pressure reaction of this matter as it cools and becomes more compact. The rebound from the sharp pressure increase at the inner region of collapse thus creates a shock for infalling envelope material some 50–100 km above the proto-neutron star surface. Infalling nuclei are decomposed in the enhanced density region of this shock and from intense neutrino interactions, thus further consuming nuclear binding energy as infalling heavy nuclei are converted to $\alpha$ particles and free nucleons. A nuclear statistical equilibrium establishes in this region, while infalling nuclei have become neutron rich from electron captures. Energies are sufficient so that nuclear reactions can shift this equilibrium to include 56Ni as matter expands and cools [79]. If a supernova explosion can be launched, expansion of the central region then rapidly reduces temperatures and densities, thus affecting nuclear reactions and cooling properties of the materials during explosion. The nuclear statistical equilibrium evolves into freeze-out with increased abundance of $\alpha$ particles. The outward-propagating shock is expected to lead to explosive nucleosynthesis as it runs through the envelope of the surrounding star that has had no time to fall in [464]. Explosive nucleosynthesis is defined as the network of nuclear reactions occurring in the envelope as an outward moving shock heats this matter to nuclear burning temperatures. During expansion, nuclear reactions fade and freeze out characteristic abundances that reflect nucleosynthesis in central regions and the expansion and cooling behavior of matter [see 295, 296, for pre-supernova and explosive nuclear processing]. The nucleosynthesis products from this inner region characteristically reflect those conditions [see 315, for a case study; such measurements are addressed in section 4.1].
@@
The special regions near the proto-neutron star, where electron captures and inverse $\beta$ decays neutronize the nuclear composition, had also been considered as plausible locations for the rapid neutron capture (r-) process [528] to occur. The early idea of r process nucleosynthesis in the high-entropy wind behind the expanding supernova shock [140] now appears unrealistic, as recent simulations of gravitational collapses rather suggest a proton-rich composition here , i.e., $Y_{e}\geq 0.5$ [497, 88]. Consequently, this has led to the suggestion of a p process occurring in this region. On the other hand, the plasma flows near a magnetically-confined jet may also create the required low-$Y_{e}$ conditions for a successful r process in magnetized gravitational-collapse variants. Also, a neutrino-related p process is being discussed for this inner supernova region. In this context it is interesting that oscillations of neutrinos in the inner supernova may be relevant and affect conditions for heavy-element synthesis [348]. The assessment of impacts on nucleosynthesis is difficult: neutrino transport and their nuclear interactions require sophisticated modeling [see, e.g. 230, 438, 145, for details].
@@ 3.2.2 Stars in binary systems
Supernovae from white dwarfs in binary systems, which result from evolution of more-massive stars that went through helium burning and thus produce a carbon-oxygen degenerate remnant occur as an ignition of central carbon fusion. The steep reaction-rate rise with temperature for this reaction leads to a runaway, as the nuclear energy release at densities near 1013g cm-3 does not immediately expand and cool the burning region, due to the degeneracy of white-dwarf matter. This results in disruption the entire white dwarf. These are called supernovae type Ia, or thermonuclear supernovae [see 304, 410, for recent reviews].
@@
Any explosive burning-expansion dynamics leaves its imprints, as the nuclear reaction flow freezes out: major expected nucleosynthesis products are 56Ni from central regions, and incompletely-burned products or unburnt carbon and oxygen from further-out regions [see 430, for a review of SN Ia nucleosynthesis]. Intermediate-mass nuclei are expected to show a characteristic abundance distribution with peaks for $\alpha$ multiples. The ratio of intermediate-mass isotopes to those of the iron group provides a diagnostic of the completeness of nuclear burning across the white dwarf [326].
@@
Accumulation of hydrogen from a companion may occur in different ways: Accreting at a small-enough rate of 10-8 to 10-10 M, hydrogen ignition already during accretion can be avoided (see below for the resulting nova phenomenon). At somewhat higher accretion rates, hydrogen burning during accretion would occur slow enough to end up as steady-state burning during accretion, and lead to the phenomenon of the supersoft X-ray sources [246, 384] as hydrogen is converted to helium. Now helium would end up to be deposited on the surface of the white dwarf. The ignition of helium burning (the 3-$\alpha$ process, see section 2) is expected above a critical mass of order 10-4 M. It would occur in a flash, sending a shock into the underlying white dwarf star. This shock may be sufficient to trigger the nuclear carbon fusion reaction in the central regions of the white dwarf, thus again leading to the thermonuclear explosion of the entire white dwarf. This is one of the sub-Chandrasekhar scenarios discussed to trigger supernovae of type Ia, complementing the above-discussed Chandrasekhar limiting mass scenario. Other such triggering precursors of thermonuclear supernovae could be accretion of helium from a companion star that already has lost its hydrogen envelope, or the more or less-violent collision of two white dwarfs. In each such scenario, the nuclear ignition and flash of helium burning provides the initial trigger for the explosion of the white dwarf, launching its violent carbon fusion reactions in central regions. It is a current topic of research to identify the relevance and detailed evolution of each of these evolutionary pathways [see, e.g. 304, for a detailed recent discussion].
@@
For neutron stars, a binary companion star and its mass transfer also may trigger explosive nuclear burning. But due to the large gravitational binding, it is difficult to release sufficient energy to liberate neutron star matter or its ashes against these $\simeq$ 200 MeV per proton of gravitational energy. Although some modest material release (of order 0.1-1% of the entire neutron-star envelope, i.e. 10-14M) has been predicted [204, 530] through, e.g., radiation-driven winds, such events are probably irrelevant for enriching interstellar gas with nucleosynthesis products. But the resulting type-I X-ray bursts observationally signify explosions on or near the the surface of a neutron star [see 165, for a review]. High densities of the gravitationally-compressed matter enable proton and $\alpha$ capture reactions to proceed beyond the hot CNO cycle reactions, and further hydrogen-burning reactions proceed along the nuclear chart up to Sn in the rapid proton capture (rp) process [423, 530, 146, 243]. Ignition of remaining ashes from this rp process, and in particular abundant carbon ashes somewhat deeper in the crust result in superbursts, more vigorous as the nuclear-energy liberation rate is higher. The light curve of such bursts is characterized by an initial rise that reflects ignition of nuclear burning, expansion and thus cooling of the burning site stops the rise, giving way to a gradual decline that is characterized by the radioactive afterglow of the mix of nuclei that resulted from nuclear burning [220]. Complexities arise, e.g., from neutron star spin [167], and from cooling properties including neutrino emission [330], and interfere with a straightforward interpretation of the lightcurve tail in terms of radioactive content of the burning ashes [181]. Nevertheless, measuring lightcurves and their variations at high time resolution provides information on the explosive burning, and details from oscillations can be related to material dynamics during the bursts [502].
@@ 3.3 Extremes of matter and energy
A prominent example is the Big Bang: Here, during expansive cooling, there is a time window where particle collisions which are sufficiently energetic for nuclear reactions are frequent, so that nuclear transformations lead to compositional changes. This period of Big Bang nucleosynthesis fades as matter expands, reducing collision frequencies and relative particle energies. Nucleosynthesis calculations for this case rely on a network of reactions among light nuclei, which mostly have been determined experimentally in underground laboratories [84] Therefore, the agreement of big bang nucleosynthesis with observations of primordial gas abundances of 4He, 3He, D, and 7Li at a baryon-to-photon ratio as inferred from measurement of the cosmic microwave background radiation by the Planck satellite is one of the pilars of the Big Bang model and the commonly-agreed standard model of cosmology [see 84, for a review of Big Bang nucleosynthesis]. The tension reported earlier for 7Li, the cosmological Li problem, has recently been interpreted as due to systematics in observations of 7Li in stars: atmospheric processes lead to mixing of the outer layers of these stars, and selection of different subsets of stars revealed systematic effects that can explain observational abundance biases [262, 341]. In the Big Bang, by contrast to explosions of compact stars, the total energy density is dominated by the initial thermal energy, i.e., by the photon content of the early universe [see 91, 83, 380, 191, for recent reviews]. The initial material composition is just composed of protons and neutrons after freeze-out of neutrino reactions, and the lack of stable nuclei of masses 5 and 8 provides a bottleneck for nuclear reactions producing heavier elements, and beyond the synthesis of helium not much has time to happen before temperature and density freeze out nuclear reactions [380]. There are, however, discussions on variants of the early evolution of the universe. One of the ideas relevant to nucleosynthesis is the suggestion of an inhomogeneous big bang that also could produce heavy elements with an r-process pattern, thus solving the issues related to determination of the source of the surprisingly homogeneous r-process ingestion from the early to the current universe [see 406, 396, for two alternative scenarios].
@@
In a dynamical situation such as gravitational collapse, the properties of the inner compressed matter in the newly-forming neutron star has an important role for the progress of collapse, and its eventual reversal into an explosion. Compressibility of matter beyond nuclear densities in neutron stars is characterized by the equation of state. This describes the pressure increase with increasing compression of matter [288]. Herein, the repulsive characteristics of the strong nuclear force are decisive at the extreme end of high densities that are reached in gravitational collapse, and nuclear phase space aspects in less extreme parts of the parameter space. Measurements of compressibility have indirectly been inferred from the merging of neutron stars in kilonovae [see 332, for a review]. Although few events only have been measured, the gravitational radiation signature allows an analysis of compressibility [see 288, for its discussion]. On the other hand, measurements on heavy nuclei in the laboratory provide an independent check and assessment of nuclear-matter properties, such as through measurements of heavy-ion collision aspects and giant resonances [170].
@@ 4 Nuclear astronomy
Abundances are best determined for the Sun. The result shown in Figure 9 for our solar system is determined after combining spectroscopy of solar radiation [25, 27] with isotopic analyses of meteoritic samples [307, 305]. (For a discussion of the methods and accuracy of stellar photospheric abundance determinations see reviews in [13, 234], and an assessment of solar abundances from photosphere, meteoritics, solar wind, and helioseismology [42]. It is unlikely to obtain such a precise measurement for other regions even in the nearby universe. But a multitude of messengers can be exploited to study nuclear physics in cosmic environments. Nevertheless, this pattern of elemental abundances appears to be largely the same throughout the current universe and its star-forming galaxies [see 136, for $\alpha$ elements in star-forming regions] [see 327, for coupling of Galactic abundances] [see 444, for a review of universality of the r process pattern]. Even galaxies of the early universe with low metallicities show surprisingly similar nucleosynthesis signatures of the processes discussed above , although a special role for first-generation stars appears prominent [408]. Note that different star formation histories and galaxy collisions, however, make such analyses uncertain, and an evolution of different enrichment histories for the various nucleosynthesis sources is likely, and part of chemical-evolution modeling (see Cumulative Nucleosynthesis below).
@@ 4.1 Nucleosynthesis events and their ejecta
Figure 11: Left: The combined spectrum of neutrino luminosities shows different components, prominently nuclear components from the Sun with 7Be and CNO reaction contributions. Right: The CNO component of solar neutrinos as measured with Borexino. The red line shows the CNO neutrino component as fitted in the analysis, the purple line marks the total fit including all background types. The best signal to noise is obtained in the yellow-marked region (see text). (Updated by [19] from [50]).
@@
Neutrinos are a direct product of nuclear reactions, with characteristic neutrino spectra reflecting specific reactions. Due to their penetrability, they escape from the dense nuclear reaction site even through massive envelopes, and thus can reach a remote observer. Cosmic sources are expected to create neutrinos through a variety of processes, each with a characteristic energy spectrum. Shown in Figure 11 (left) are the cosmological neutrinos as they decoupled from the plasma of the early universe (lowest energies), then the solar neutrinos from nuclear reactions (MeV energies). Towards higher energies, supernova neutrinos from the cooling proto-neutron star with a continuum spectrum centered at 10–30 MeV is expected from theory, consistent with the few events recorded from SN1987A. Hadronic interactions of cosmic rays in the Earth atmosphere form powerlaw-like spectra towards higher energies that have been measured in detail by, e.g., IceCube and Kamiokande experiments. Probably the jets created in black hole formation or accretion such as in gamma-ray bursts and active galactic nuclei, respectively, will create similar neutrino spectra at even higher energies, corresponding to the most-energetic cosmic rays.
@@
For the Sun, characteristic energies of neutrinos from nuclear reactions are around the MeV scale, emitted by reactions along the hydrogen burning sequences [390, 7]. These are dominated by neutrinos from the pp chain reactions, which are mostly of energies below 0.5 MeV and therefore hard to detect, as neutrino cross sections steeply rise with energies. The pep neutrinos from the initial deuterium formation in the pp chain however are mono-energetic at 1.44 MeV, and thus a target of measurements. Solar neutrinos include CNO cycle neutrinos with energies up to 1.7 MeV, from $\beta$ decays of 13N, 15O, and 17F, and also 7B neutrinos with two lines at energies 0.86 and 0.36 MeV, that are most intense and thus easier to detect [150]. This is a challenging energy range, because many processes of nuclear physics in the detector and its surroundings lead to similar signals, and it is difficult to recognize such background against the few desired signatures. Moreover, spectroscopic resolution at the 0.1 MeV level or better is required to attribute the neutrinos to specific nuclear reactions. The discovery of neutrinos from the Sun and assessment of its measurement in terms of modeling the Sun and understanding the nature of neutrinos has been a major triumph of nuclear astrophysics. An underground experiment based on conversion of chlorine to argon by neutrinos reported a flux limit that lay a factor 7 below the flux of (the most energetic) 8B neutrinos predicted by models of the Sun [96] [also see the corresponding Nobel prize lecture by 95]). The SAGE and GALLEX experiments similarly used chemical extraction methods to measure solar neutrinos, and confirmed the data recorded by the Homestad experiment. But more convincing appeared the direct neutrino event detection in the Kamiokande imaging water Cherenkov experiment, which later verified this discrepancy and established the solar neutrino problem, at the same time discussing its potential solution through neutrino flavor oscillations [265]. The heavy-water based SNO experiment operated in Canada at the Sudbury Neutrino Laboratory then added the capability to measure the neutral-current neutrino reactions through weak-interaction neutron release from deuterium (producing a characteristic neutron capture $\gamma$-ray line at 6.25 MeV), in addition to their measurements of the charged-current $\nu_{e}$ interactions, thus providing experimental proof of neutrino flavor oscillations [10]. The theoretical predictions and experimental confirmations of neutrino flavor oscillations were assessed as the explanation of the measured deficit: electron neutrinos emitted in the nuclear reactions change flavor along their trajectories, so that a lower electron neutrino flux interacts with detectors on Earth. Including these neutrino properties, current solar-model predictions now are in general agreement with measured values [214].
@@
Also for thermonuclear supernovae, a first successful case of observations of nuclear gamma rays on nearby SN2014J turns out to be quite informative. Even though it is expected that centrally-ignited carbon burning and its main radioactive product of 56Ni should be occulted by the massive white-dwarf envelope for about three months, characteristic $\gamma$ rays from 56Ni decay (with a radioactive lifetime $\tau$ of 9 days) at 158 and 812 keV could be seen [112]. This illustrates that either the envelope and central burning ashes adopt a more-complex morphology, or else some 56Ni had been synthesized near the white-dwarf surface, possibly from helium accreted from a companion, in a non-Chandrasekhar scenario (explained in section 3.2 on Stellar Explosions) [see also discussion in 112]. From this same SN2014J event, the characteristic $\gamma$-ray lines from the decay of 56Ni-decay daughter 56Co (with radioactive lifetime $\tau$ of 111 days) were measured by both INTEGRAL gamma-ray telescopes [76]. They appear characteristically Doppler broadened from the explosion. But beyond, SPI $\gamma$-ray spectra in high-resolution detail show indications of a non-spherical explosion, i.e. individual bulk-velocity spikes rather than a broad and smooth Doppler broadening at all times [113]. Observations of supernova light is likely to diffuse and probably not show such information, as bolometric absorption and down-scattering in energy within a massive envelope is required for such radiation conversion.
Other radioactive by-products of nucleosynthesis might be detected in their characteristic $\gamma$-ray lines [e.g. 97, for science perspectives with a typical next-generation telescope]. Candidates are 22Na (radioactive lifetime $\tau\simeq$3.8 y), expected to be produced in novae [240] or in massive-star core collapses [526, 379], or 126Sn (radioactive lifetime $\tau\simeq$3.3 105 y), predicted to be produced in kilonovae [303]. For novae, none of the candidate lines from positron annihilation of shortlived $\beta^{+}$ nuclei (511 keV line), 7Be, or 22Na, could be detected [436]. An interesting tentative detection of 7Be from INTEGRAL data [225] appears promising for future instruments.

### v2

@@ 3.2.1 Single stars
As a result of these various fading central energy sources towards the end of stellar evolution of stars with masses exceeding $\simeq$ 8 M, gravitational collapse will take over, and proceed faster with electron capture and hence neutronization increasing in central regions [see 232, for a recent review]. This neutronization of matter near and in the newly-forming neutron star in the center of gravitational collapse results in an additional and very intense burst of electron neutrinos. This is one of the significant messengers for astronomical measurement (see section 4 on Nuclear Astronomy).
The collapse is halted when central density reaches values above nuclear saturation density beyond 2.7 1014g cm-3, as compression then faces the repulsive part of the strong nuclear potential. This is reflected in the equation of state of matter in the newly-forming neutron star [230], which describes the pressure reaction of this matter as it cools and becomes more compact. The rebound from the sharp pressure increase at the inner region of collapse thus creates a shock for infalling envelope material some 50–100 km above the proto-neutron star surface. Infalling nuclei are decomposed in the enhanced density region of this shock and from intense neutrino interactions, thus further consuming nuclear binding energy as infalling heavy nuclei are converted to $\alpha$ particles and free nucleons. A nuclear statistical equilibrium establishes in this region, while infalling nuclei have become neutron rich from electron captures. Energies are sufficient so that nuclear reactions can shift this equilibrium to include 56Ni as matter expands and cools [80]. If a supernova explosion can be launched, expansion of the central region then rapidly reduces temperatures and densities, thus affecting nuclear reactions and cooling properties of the materials during explosion. The nuclear statistical equilibrium evolves into freeze-out with increased abundance of $\alpha$ particles. The outward-propagating shock is expected to lead to explosive nucleosynthesis as it runs through the envelope of the surrounding star that has had no time to fall in [468]. Explosive nucleosynthesis is defined as the network of nuclear reactions occurring in the envelope as an outward moving shock heats this matter to nuclear burning temperatures. During expansion, nuclear reactions fade and freeze out characteristic abundances that reflect nucleosynthesis in central regions and the expansion and cooling behavior of matter [see 296, 297, for pre-supernova and explosive nuclear processing]. The nucleosynthesis products from this inner region characteristically reflect those conditions [see 317, for a case study; such measurements are addressed in section 4.1].
@@
The special regions near the proto-neutron star, where electron captures and inverse $\beta$ decays neutronize the nuclear composition, had also been considered as plausible locations for the rapid neutron capture (r-) process [532] to occur. The early idea of r process nucleosynthesis in the high-entropy wind behind the expanding supernova shock [141] now appears unrealistic, as recent simulations of gravitational collapses rather suggest a proton-rich composition here , i.e., $Y_{e}\geq 0.5$ [501, 89]. Consequently, this has led to the suggestion of a p process occurring in this region. On the other hand, the plasma flows near a magnetically-confined jet may also create the required low-$Y_{e}$ conditions for a successful r process in magnetized gravitational-collapse variants. Also, a neutrino-related p process is being discussed for this inner supernova region. In this context it is interesting that oscillations of neutrinos in the inner supernova may be relevant and affect conditions for heavy-element synthesis [350]. The assessment of impacts on nucleosynthesis is difficult: neutrino transport and their nuclear interactions require sophisticated modeling [see, e.g. 231, 441, 146, for details].
@@ 3.2.2 Stars in binary systems
Supernovae from white dwarfs in binary systems, which result from evolution of more-massive stars that went through helium burning and thus produce a carbon-oxygen degenerate remnant occur as an ignition of central carbon fusion. The steep reaction-rate rise with temperature for this reaction leads to a runaway, as the nuclear energy release at densities near 1013g cm-3 does not immediately expand and cool the burning region, due to the degeneracy of white-dwarf matter. This results in disruption the entire white dwarf. These are called supernovae type Ia, or thermonuclear supernovae [see 305, 413, for recent reviews].
@@
Any explosive burning-expansion dynamics leaves its imprints, as the nuclear reaction flow freezes out: major expected nucleosynthesis products are 56Ni from central regions, and incompletely-burned products or unburnt carbon and oxygen from further-out regions [see 433, for a review of SN Ia nucleosynthesis]. Intermediate-mass nuclei are expected to show a characteristic abundance distribution with peaks for $\alpha$ multiples. The ratio of intermediate-mass isotopes to those of the iron group provides a diagnostic of the completeness of nuclear burning across the white dwarf [328].
@@
Accumulation of hydrogen from a companion may occur in different ways: Accreting at a small-enough rate of 10-8 to 10-10 M, hydrogen ignition already during accretion can be avoided (see below for the resulting nova phenomenon). At somewhat higher accretion rates, hydrogen burning during accretion would occur slow enough to end up as steady-state burning during accretion, and lead to the phenomenon of the supersoft X-ray sources [247, 387] as hydrogen is converted to helium. Now helium would end up to be deposited on the surface of the white dwarf. The ignition of helium burning (the 3-$\alpha$ process, see section 2) is expected above a critical mass of order 10-4 M. It would occur in a flash, sending a shock into the underlying white dwarf star. This shock may be sufficient to trigger the nuclear carbon fusion reaction in the central regions of the white dwarf, thus again leading to the thermonuclear explosion of the entire white dwarf. This is one of the sub-Chandrasekhar scenarios discussed to trigger supernovae of type Ia, complementing the above-discussed Chandrasekhar limiting mass scenario. Other such triggering precursors of thermonuclear supernovae could be accretion of helium from a companion star that already has lost its hydrogen envelope, or the more or less-violent collision of two white dwarfs. In each such scenario, the nuclear ignition and flash of helium burning provides the initial trigger for the explosion of the white dwarf, launching its violent carbon fusion reactions in central regions. It is a current topic of research to identify the relevance and detailed evolution of each of these evolutionary pathways [see, e.g. 305, for a detailed recent discussion].
@@
For neutron stars, a binary companion star and its mass transfer also may trigger explosive nuclear burning. But due to the large gravitational binding, it is difficult to release sufficient energy to liberate neutron star matter or its ashes against these $\simeq$ 200 MeV per proton of gravitational energy. Although some modest material release (of order 0.1-1% of the entire neutron-star envelope, i.e. 10-14M) has been predicted [205, 534] through, e.g., radiation-driven winds, such events are probably irrelevant for enriching interstellar gas with nucleosynthesis products. But the resulting type-I X-ray bursts observationally signify explosions on or near the the surface of a neutron star [see 166, for a review]. High densities of the gravitationally-compressed matter enable proton and $\alpha$ capture reactions to proceed beyond the hot CNO cycle reactions, and further hydrogen-burning reactions proceed along the nuclear chart up to Sn in the rapid proton capture (rp) process [426, 534, 147, 244]. Ignition of remaining ashes from this rp process, and in particular abundant carbon ashes somewhat deeper in the crust result in superbursts, more vigorous as the nuclear-energy liberation rate is higher. The light curve of such bursts is characterized by an initial rise that reflects ignition of nuclear burning, expansion and thus cooling of the burning site stops the rise, giving way to a gradual decline that is characterized by the radioactive afterglow of the mix of nuclei that resulted from nuclear burning [221]. Complexities arise, e.g., from neutron star spin [168], and from cooling properties including neutrino emission [332], and interfere with a straightforward interpretation of the lightcurve tail in terms of radioactive content of the burning ashes [182]. Nevertheless, measuring lightcurves and their variations at high time resolution provides information on the explosive burning, and details from oscillations can be related to material dynamics during the bursts [506].
@@ 3.3 Extremes of matter and energy
A prominent example is the Big Bang: Here, during expansive cooling, there is a time window where particle collisions which are sufficiently energetic for nuclear reactions are frequent, so that nuclear transformations lead to compositional changes. This period of Big Bang nucleosynthesis fades as matter expands, reducing collision frequencies and relative particle energies. Nucleosynthesis calculations for this case rely on a network of reactions among light nuclei, which mostly have been determined experimentally in underground laboratories [85] Therefore, the agreement of big bang nucleosynthesis with observations of primordial gas abundances of 4He, 3He, D, and 7Li at a baryon-to-photon ratio as inferred from measurement of the cosmic microwave background radiation by the Planck satellite is one of the pilars of the Big Bang model and the commonly-agreed standard model of cosmology [see 85, for a review of Big Bang nucleosynthesis]. The tension reported earlier for 7Li, the cosmological Li problem, has recently been interpreted as due to systematics in observations of 7Li in stars: atmospheric processes lead to mixing of the outer layers of these stars, and selection of different subsets of stars revealed systematic effects that can explain observational abundance biases [263, 343]. In the Big Bang, by contrast to explosions of compact stars, the total energy density is dominated by the initial thermal energy, i.e., by the photon content of the early universe [see 92, 84, 383, 192, for recent reviews]. The initial material composition is just composed of protons and neutrons after freeze-out of neutrino reactions, and the lack of stable nuclei of masses 5 and 8 provides a bottleneck for nuclear reactions producing heavier elements, and beyond the synthesis of helium not much has time to happen before temperature and density freeze out nuclear reactions [383]. There are, however, discussions on variants of the early evolution of the universe. One of the ideas relevant to nucleosynthesis is the suggestion of an inhomogeneous big bang that also could produce heavy elements with an r-process pattern, thus solving the issues related to determination of the source of the surprisingly homogeneous r-process ingestion from the early to the current universe [see 409, 399, for two alternative scenarios].
@@
In a dynamical situation such as gravitational collapse, the properties of the inner compressed matter in the newly-forming neutron star has an important role for the progress of collapse, and its eventual reversal into an explosion. Compressibility of matter beyond nuclear densities in neutron stars is characterized by the equation of state. This describes the pressure increase with increasing compression of matter [289]. Herein, the repulsive characteristics of the strong nuclear force are decisive at the extreme end of high densities that are reached in gravitational collapse, and nuclear phase space aspects in less extreme parts of the parameter space. Measurements of compressibility have indirectly been inferred from the merging of neutron stars in kilonovae [see 334, for a review]. Although few events only have been measured, the gravitational radiation signature allows an analysis of compressibility [see 289, for its discussion]. On the other hand, measurements on heavy nuclei in the laboratory provide an independent check and assessment of nuclear-matter properties, such as through measurements of heavy-ion collision aspects and giant resonances [171].
@@ 4 Nuclear astronomy
Abundances are best determined for the Sun. The result shown in Figure 9 for our solar system is determined after combining spectroscopy of solar radiation [26, 28] with isotopic analyses of meteoritic samples [308, 306]. (For a discussion of the methods and accuracy of stellar photospheric abundance determinations see reviews in [14, 235], and an assessment of solar abundances from photosphere, meteoritics, solar wind, and helioseismology [43]. It is unlikely to obtain such a precise measurement for other regions even in the nearby universe. But a multitude of messengers can be exploited to study nuclear physics in cosmic environments. Nevertheless, this pattern of elemental abundances appears to be largely the same throughout the current universe and its star-forming galaxies [see 137, for $\alpha$ elements in star-forming regions] [see 329, for coupling of Galactic abundances] [see 447, for a review of universality of the r process pattern]. Even galaxies of the early universe with low metallicities show surprisingly similar nucleosynthesis signatures of the processes discussed above , although a special role for first-generation stars appears prominent [411]. Note that different star formation histories and galaxy collisions, however, make such analyses uncertain, and an evolution of different enrichment histories for the various nucleosynthesis sources is likely, and part of chemical-evolution modeling (see Cumulative Nucleosynthesis below).
@@ 4.1 Nucleosynthesis events and their ejecta
Figure 11: Top: The combined spectrum of neutrino luminosities from the variety of sources shows all different components, with the prominent nuclear components from the Sun, with its dominating proton-chain continua and the 7Be and 8B lines, with continua from the CNO reactions adding spectral modulation [from 493]. Bottom: The CNO component of solar neutrinos as measured with Borexino. The red line shows the CNO neutrino component as fitted in the analysis, the purple line marks the total fit including all background types. The best signal to noise is obtained in the yellow-marked region (see text). (Updated by [20] from [51]).
@@
Neutrinos are a direct product of nuclear reactions, with characteristic neutrino spectra reflecting specific reactions. Due to their penetrability, they escape from the dense nuclear reaction site even through massive envelopes, and thus can reach a remote observer. Cosmic sources are expected to create neutrinos through a variety of processes, each with a characteristic energy spectrum. Shown in Figure 11 (left) are the cosmological neutrinos as they decoupled from the plasma of the early universe (lowest energies), then the solar neutrinos from nuclear reactions (MeV energies) [493, 215]. Towards higher energies, supernova neutrinos from the cooling proto-neutron star with a continuum spectrum centered at 10–30 MeV is expected from theory, consistent with the few events recorded from SN1987A [309]. Hadronic interactions of cosmic rays in the Earth atmosphere form powerlaw-like spectra towards higher energies that have been measured in detail by, e.g., IceCube and Kamiokande experiments [1, 466]. Probably the jets created in black hole formation or accretion such as in gamma-ray bursts and active galactic nuclei, respectively, will create similar neutrino spectra at even higher energies, corresponding to the most-energetic cosmic rays [355].
@@
For the Sun, characteristic energies of neutrinos from nuclear reactions are around the MeV scale, emitted by reactions along the hydrogen burning sequences [393, 8]. These are dominated by neutrinos from the pp chain reactions, which are mostly of energies below 0.5 MeV and therefore hard to detect, as neutrino cross sections steeply rise with energies. The pep neutrinos from the initial deuterium formation in the pp chain however are mono-energetic at 1.44 MeV, and thus a target of measurements. Solar neutrinos include CNO cycle neutrinos with energies up to 1.7 MeV, from $\beta$ decays of 13N, 15O, and 17F, and also 7B neutrinos with two lines at energies 0.86 and 0.36 MeV, that are most intense and thus easier to detect [151]. This is a challenging energy range, because many processes of nuclear physics in the detector and its surroundings lead to similar signals, and it is difficult to recognize such background against the few desired signatures. Moreover, spectroscopic resolution at the 0.1 MeV level or better is required to attribute the neutrinos to specific nuclear reactions. The discovery of neutrinos from the Sun and assessment of its measurement in terms of modeling the Sun and understanding the nature of neutrinos has been a major triumph of nuclear astrophysics. An underground experiment based on conversion of chlorine to argon by neutrinos reported a flux limit that lay a factor 7 below the flux of (the most energetic) 8B neutrinos predicted by models of the Sun [97] [also see the corresponding Nobel prize lecture by 96]). The SAGE and GALLEX experiments similarly used chemical extraction methods to measure solar neutrinos, and confirmed the data recorded by the Homestad experiment. But more convincing appeared the direct neutrino event detection in the Kamiokande imaging water Cherenkov experiment, which later verified this discrepancy and established the solar neutrino problem, at the same time discussing its potential solution through neutrino flavor oscillations [266]. The heavy-water based SNO experiment operated in Canada at the Sudbury Neutrino Laboratory then added the capability to measure the neutral-current neutrino reactions through weak-interaction neutron release from deuterium (producing a characteristic neutron capture $\gamma$-ray line at 6.25 MeV), in addition to their measurements of the charged-current $\nu_{e}$ interactions, thus providing experimental proof of neutrino flavor oscillations [11]. The theoretical predictions and experimental confirmations of neutrino flavor oscillations were assessed as the explanation of the measured deficit: electron neutrinos emitted in the nuclear reactions change flavor along their trajectories, so that a lower electron neutrino flux interacts with detectors on Earth. Including these neutrino properties, current solar-model predictions now are in general agreement with measured values [215].
@@
Also for thermonuclear supernovae, a first successful case of observations of nuclear gamma rays on nearby SN2014J turns out to be quite informative. Even though it is expected that centrally-ignited carbon burning and its main radioactive product of 56Ni should be occulted by the massive white-dwarf envelope for about three months, characteristic $\gamma$ rays from 56Ni decay (with a radioactive lifetime $\tau$ of 9 days) at 158 and 812 keV could be seen [113]. This illustrates that either the envelope and central burning ashes adopt a more-complex morphology, or else some 56Ni had been synthesized near the white-dwarf surface, possibly from helium accreted from a companion, in a non-Chandrasekhar scenario (explained in section 3.2 on Stellar Explosions) [see also discussion in 113]. From this same SN2014J event, the characteristic $\gamma$-ray lines from the decay of 56Ni-decay daughter 56Co (with radioactive lifetime $\tau$ of 111 days) were measured by both INTEGRAL gamma-ray telescopes [77]. They appear characteristically Doppler broadened from the explosion. But beyond, SPI $\gamma$-ray spectra in high-resolution detail show indications of a non-spherical explosion, i.e. individual bulk-velocity spikes rather than a broad and smooth Doppler broadening at all times [114]. Observations of supernova light is likely to diffuse and probably not show such information, as bolometric absorption and down-scattering in energy within a massive envelope is required for such radiation conversion.
Other radioactive by-products of nucleosynthesis might be detected in their characteristic $\gamma$-ray lines [e.g. 98, for science perspectives with a typical next-generation telescope]. Candidates are 22Na (radioactive lifetime $\tau\simeq$3.8 y), expected to be produced in novae [241] or in massive-star core collapses [530, 382], or 126Sn (radioactive lifetime $\tau\simeq$3.3 105 y), predicted to be produced in kilonovae [304]. For novae, none of the candidate lines from positron annihilation of shortlived $\beta^{+}$ nuclei (511 keV line), 7Be, or 22Na, could be detected [439]. An interesting tentative detection of 7Be from INTEGRAL data [226] appears promising for future instruments.